Feauto Manufacturing Corporation has a traditional costing system in which it applies manufacturing overhead to its products using a predetermined overhead rate based on direct labor-hours (DLHs). The company has two products, I63E and E76I, about which it has provided the following data: I63E E76I Direct materials per unit $ 21.10 $ 63.30 Direct labor per unit $ 18.00 $ 54.00 Direct labor-hours per unit 0.90 2.70 Annual production (units) 54,000 18,000 The company's estimated total manufacturing overhead for the year is $2,608,200 and the company's estimated total direct labor-hours for the year is 97,200. The company is considering using a form of activity-based costing to determine its unit product costs for external reports. Data for this proposed activity-based costing system appear below: Activities and Activity Measures Estimated Overhead Cost Assembling products (DLHs) $ 1,458,000 Preparing batches (batches) 178,200 Product support (product variations) 972,000 Total $ 2,608,200 Expected Activity I63E E76I Total DLHs 48,600 48,600 97,200 Batches 1,240 670 1,910 Product variations 2,480

1,340 3,820
Answer #1

Estimaated labour hour = (54000*.9+18000*2.7) = 97200 Hours

Overhead rate = 2,608,200/97200 = 26.83 per hour

The manufacturing overhead that would be applied to a unit of product I63E under the company's traditional costing system is closest to = 26.83*.90 = 24 per unit